# Gravity and fluid dynamic correspondence on a null hypersurface: inconsistencies and advancement

### v1

@@ 2.1 Relativistic fluid and its non-relativistic counterpart
Moreover, the components of the energy-momentum tensor $T^{00}_{(fluid)}$, $T^{0\alpha}_{(fluid)}$ and $T^{\alpha\beta}_{(fluid)}$ represent energy density, momentum density and stress respectively. Similarly, the baryon current is also conserved. More importantly, there are five independent parameters, three components of fluid velocity, pressure, and energy density, that are required to describe fluid completely. In order to uniquely determine these unknown parameters, five independent equations are required. These equations can be given as follows. It is known that if one takes a timelike projection of the conservation law 1, i.e.
$v_{b}\nabla_{a}T^{ab}_{(fluid)}=0~{},$ (2)
@@
it yields the energy conservation relation of the fluid, where $v_{b}$ is the four-velocity of the fluid. On the other hand, if one considers the spacelike projection of 1, i.e.
$\Delta^{c}_{b}~{}\nabla_{a}T^{ab}_{(fluid)}=0~{},$ (3)
@@
provides the momentum conservation of the fluid, where $\Delta^{a}_{b}=\delta^{a}_{b}+v^{a}v_{b}$ is the projection tensor, which is orthogonal to the four-velocity $v_{a}$. Note that the above 3 is a set of three equations. In addition to that, there is a continuity equation, coming from the baryon current conservation, which at the non-relativistic limit, is given as
$\frac{\partial\rho}{\partial t}+\mathbf{\nabla}\cdot(\rho\mathbf{v})=0~{},$ (4)
where $\rho$ is the volumetric density of the fluid. Thus, together, 2, 3 and 4 make a set of five equations that are required to determine the five unknown quantities stated earlier.
$D^{N}\rho+\rho\theta^{N}=0~{},$
$\rho D^{N}\epsilon^{N}+P\theta^{N}+\partial_{\mu}Q^{\mu}-\zeta(\theta^{N})^{2}-2\eta(\sigma^{N})^{\mu\nu}\sigma^{N}_{\mu\nu}=0~{},$
$\partial^{\mu}P+\rho D^{N}v^{\mu}-\partial_{\nu}\Big{[}\zeta\theta^{N}\delta^{\mu\nu}+2\eta(\sigma^{N})^{\mu\nu}\Big{]}=0~{},$ (5)
where
$D^{N}\equiv\partial_{t}+v^{\mu}\partial_{\mu}~{},$
$\sigma^{N}_{\mu\nu}\equiv\partial_{(\mu}v_{\nu)}-\frac{1}{3}\delta_{\mu\nu}\theta^{N}~{},$
$\theta^{N}=\partial_{\mu}v^{\mu}~{}.$ (6)
Here the index ``$N"$ is used for Navier-Stokes (non-relativistic) fluid, which is used to distinguish the physical quantities from the gravitational counterpart. If one uses the continuity equation in the energy conservation and momentum conservation relation, one obtains the modified form of energy and momentum conservation relations, where the energy conservation relation is given as
$D^{N}(\rho\epsilon^{N})+\Big{(}\rho\epsilon^{N}+P\Big{)}\theta^{N}+\partial_{\mu}Q^{\mu}-\zeta(\theta^{N})^{2}-2\eta(\sigma^{N})^{\mu\nu}\sigma^{N}_{\mu\nu}=0~{},$ (7)
and the momentum conservation relation is given as
$D^{N}\Big{(}\rho v^{\mu}\Big{)}+\rho v^{\mu}\theta^{N}+\partial^{\mu}P-\partial_{\nu}\Big{[}\zeta\theta^{N}\delta^{\mu\nu}+2\eta(\sigma^{N})^{\mu\nu}\Big{]}=0~{}.$ (8)
8 provides the dynamics of the fluid while 7 describes its kinematics. In what follows, we show that the analogous relations of 7 and 8 exist in gravitational dynamics on a generic null hypersurface. More importantly, those relations are obtained in a conventional manner i.e. taking different projections of conservation of suitably chosen gravitational energy-momentum tensor, similar as described in 2 and 3 for usual fluid. In 3, we obtain the analogous energy-momentum tensor for the null hydrodynamics. But before that, we explain briefly the geometric construction of the null surface. This serves the purpose of realizing the geometrical quantities, defined on a generic null hypersurface, which play an important role in determining the gravitational fluid parameters. Our objective will be to consider an integrable null hypersurface $\mathcal{H}$ and then to foliate the neighborhood of the spacetime by a family of such null surfaces. We will also describe briefly the construction of an adapted coordinate system with respect to $\mathcal{H}$.
@@ 2.2 A generic null hypersurface
In the Riemannian spacetime manifold $(\mathcal{M},\textbf{g},\bm{\nabla})$, equipped with the Levi-Civita connection $\bm{\nabla}$, a null hypersurface is a surface of codimension one such the induced metric on it is degenerate. Since, in our construction, we consider a generic integrable null surface, it can be defined via a scalar function $u(x^{a})=0$. The null normal $\underline{\bm{{l}}}=l_{a}\bm{d}x^{a}$ to such a surface is given by,
@@
with $\hat{\rho}$ being a smooth scalar field defined on the null surface $\mathcal{H}$. The negative sign in Eq. (9) accounts for the fact the null generators $l^{a}$ to $\mathcal{H}$ are future-directed by an appropriate choice of the scalar function $u(x^{a})$. On account of the null nature of the normal to $\mathcal{H}$, $\bm{{l}}$ cannot be provided a unique normalization since $\bm{{l}}\cdot\bm{{l}}=0$. The fact that we have considered only a single null surface, which is the support of the null generator $\bm{{l}}$, we cannot take notions of derivatives of quantities away from the null surface $\mathcal{H}$. To circumvent this problem, following Carter [48], we consider not just a single null surface but rather a family of them. That is, we foliate the spacetime $(\mathcal{M},\textbf{g},\bm{\nabla})$ in the neighborhood of $\mathcal{H}$ by a stack of integrable null surfaces defined by $u(x^{a})=c$, where $c$ is a constant. So we have a family of integrable null surfaces $\mathcal{H}_{u}$ out of which $\mathcal{H}_{u=0}=\mathcal{H}$ is our chosen integrable null hypersurface. This construction then extends the support of the null generator $\bm{{l}}$ to not just over $\mathcal{H}$ but rather over a finite region in the spacetime. This then facilitates the notion of taking derivatives of quantities away from the null surface. It is worth mentioning that such a foliation of the ambient spacetime in the neighborhood of $\mathcal{H}$ is non-unique. However, all geometric quantities that will be described for the null surface will be independent of any such foliation that has been introduced. The integrability of the null surfaces implies that it satisfies the Frobenius identity,
@@
The fact, that the null normal lies on $\mathcal{H}$ itself, does not allow the construction of a projection tensor (that can be created out from the ambient spacetime metric and the null normal) onto the null surface. For that, we need to have the notion of a vector field that is transverse to the null hypersurface. But before introducing such a transverse vector field, let us consider a $3+1$ foliation of the null family $u(x^{a})=c$ by a stack of spacelike surfaces $\Sigma_{t}$ defined by the scalar function $t(x^{a})=\text{constant}$. The spacelike surfaces are assumed not to intersect with each other. Hence, a local neighborhood of the spacetime manifold in the vicinity of $\mathcal{H}$ can be coordinatized by $(t,x^{1},x^{2},x^{3})$, with $x^{\mu}=(x^{1},x^{2},x^{3})$ being the spatial coordinates on a given $\Sigma_{t}$. The time evolution vector field $\bm{t}=\bm{\partial_{t}}$ that connects the same spatial points $x^{\mu}$ of the neighboring slices $\Sigma_{t}$ satisfies the condition $t^{a}\partial_{a}t=1$. The cuts/intersections of the null surface $\mathcal{H}$ with the spacelike familty $\Sigma_{t}$ is defined as $\mathcal{B}^{(null)}\equiv\mathcal{H}\cap\Sigma_{t}$ which form a $2$-dimensional submanifold of $\mathcal{H}$. With this, we introduce a unique notion of a transverse auxiliary null vector field $\bm{{k}}$, defined as,
@@
where $\{\bm{e_{{A}}}\}$ are the set of two spacelike basis vectors on $\mathcal{B}^{(null)}$. The null generators $\bm{{l}}$ provide a notion of outgoing null vector field whereas the auxiliary null vector field $\bm{{k}}$ is ingoing w.r.t the codimension two submanifolds $\mathcal{B}^{(null)}$. With the null generators and the auxiliary null vector field, one can define two projection tensors
@@
On the contrary, $q^{a}_{b}$ is the projection tensor onto the two-surface $\mathcal{B}^{(null)}$ upon which both $l^{a}$ and $k^{a}$ are the normals (i.e. cross-section of $\mathcal{H}$). The properties of $q^{a}_{b}$ are as follows:
@@
Unlike $\Pi^{a}_{~{}b}$, $q^{a}_{b}$ is the induced metric on $\mathcal{B}^{(null)}$ as well. Finally, we discuss the notion of adapting a coordinate system [25] concerning our integrable null surface $\mathcal{H}$. The choice will be such that the equation describing $\mathcal{H}$ will not depend on the time parameter $t$ and will involve only the spatial coordinates $x^{\mu}$. Adapting a coordinate system with respect to $\mathcal{H}$ means that the location of the null surface has been fixed by the spatial coordinates via a judicious choice of a scalar function say $f(x^{1},x^{2},x^{2})=0$ [25]. As mentioned before, owing to the $3+1$ foliation of $\mathcal{H}$ by the family of spacelike surfaces $\Sigma_{t}$ allows us to coordinatize spacetime manifold in the vicinity of $\mathcal{H}$ by $(t,x^{1},x^{2},x^{3})$. Notice that the function $f(x^{\mu})$ is independent of the time parameter $t$. This further entails that,
$\frac{\partial u}{\partial t}\overset{\mathcal{H}}{=}0~{}.$ (15)
One appropriate choice may be to fix the location of the null surface by choosing $f(x^{1},x^{2},x^{2})=x^{1}=0$. This means that on a given $t(x^{1},x^{2},x^{3})=\text{constant}$ slice, the codimension two surface $\mathcal{B}^{(null)}$ is fixed or determined by the choice $x^{1}=0$. Then naturally, the transverse/angular coordinates on $\mathcal{B}^{(null)}$ are $x^{A}=(x^{2},x^{3})$. Along the time evolution vector field $\bm{t}=\bm{\partial_{t}}$, the spatial coordinates $(x^{2},x^{3})$ of $\mathcal{B}^{(null)}$ are fixed. Hence with this adapted choice, the coordinates parametrizing the null surface $\mathcal{H}$ are $x^{\tilde{\mu}}=(t,x^{2},x^{3})$.
It can be shown [25], that on the null surface we have, $\bm{l}\overset{\mathcal{H}}{=}\bm{t}+\bm{V}$, where $\bm{V}$ is spatial vector field that is tangent to $\mathcal{B}^{(null)}$. With respect to the adapted coordinate system, hence, we have $l^{\tilde{\mu}}\overset{\mathcal{H}}{=}(1,V^{A})$. The line element on the null surface in this adapted coordinate system can be expressed as [25, 26],
The intrinsic geometry of the null surface is provided by the first fundamental form, which is the metric $\bm{q}$ induced from the ambient metric $\bm{g}$ from the spacetime manifold $(\mathcal{M},\textbf{g},\bm{\nabla})$. The codimension of two surfaces $\mathcal{B}^{(null)}$ forms a submanifold $(\mathcal{B}^{(null)},\bm{q},\bm{~{}^{(2)}D})$, equipped with the unique metric compatible torsion-free connection $\bm{~{}^{(2)}D}$ satisfying ${}^{(2)}D_{A}q_{BC}=0$. To completely characterize the extrinsic geometry of the null surface, we need the triplet $(\theta_{ab},\Omega_{a},\kappa)$, where $\theta_{ab}$ is the second fundamental form of $\mathcal{H}$ defined as [25],
@@
In principle, the Hajicek one form is the projection of something called the rotation one-form $\omega_{a}$ onto the submanifold $(\mathcal{B}^{(null)},\bm{q},\bm{~{}^{(2)}D})$, defined as [25],
@@
Now, it could be possible that for a specific choice of the adapted coordinates, we could have the constraint that $\partial_{t}q_{AB}=0$ i.e. the induced metric $q_{AB}$ is stationary or independent of the time evolution parameter $t$. Then the second fundamental form of the null surface $\mathcal{H}$ takes exactly the same structure as that of the stress tensor of a two-dimensional viscous fluid with velocity $V^{A}$. This part will be used to relate the gravitational equations with NS fluid equations 7 and 8.
@@
$q_{a}^{~{}i}q_{b}^{~{}j}\Big{(}\pounds_{\bm{{l}}}\theta_{ab}\Big{)}=\kappa\theta_{ab}+\theta_{ai}\theta^{i}_{~{}b}-q_{a}^{~{}i}q_{b}^{~{}j}C_{minj}l^{m}l^{n}-\frac{1}{2}q_{ab}R_{ij}l^{i}l^{j}~{},$ (25)
where $C_{abcd}$ is the Weyl tensor for $(\mathcal{M},\textbf{g},\bm{\nabla})$. Taking the trace of (25) gives us the celebrated null Raychaudhuri equation (NRE),
@@
and restricting to the spatial coordinates of $(\mathcal{B}^{(null)},\bm{q},\bm{~{}^{(2)}D})$, the above (27) gives us the Hajicek equation [45]. For the case of Einstein gravity, upon analyzing the evolution equation for the Hajicek one-form in the adapted coordinate system w.r.t. $\mathcal{H}$ and using that £$\pounds_{\bm{{l}}}\Omega_{a}\overset{\mathcal{H}}{=}\pounds_{\bm{t}}\Omega_{a}+\pounds_{\bm{V}}\Omega_{a}$, we have,
This is the well-known Damour-Navier-Stokes equation [23, 50]. In fact, it was shown by Padmanabhan [26], that analyzing the DNS equation with respect to a boosted local inertial frame, the extra term $\Omega_{b}~{}^{(2)}D_{a}V^{b}$ drops out and it takes the structure of a two dimensional viscous Navier-Stokes fluid, identical to 8 for a specific choice of fluid parameters in terms of gravitational quantities (this we will elaborate later as well).
@@ 3.1 BY tensor: laying the path for null counterpart
$\delta\mathcal{A}_{WP}=\delta\mathcal{A}_{EH}+\delta\mathcal{A}_{GHY}=\delta\Big{(}\frac{1}{16\pi}\int_{\mathcal{V}}\sqrt{-g}R\ d^{4}x-\frac{\epsilon}{8\pi}\int_{\partial\mathcal{V}}\sqrt{\mathbf{h}}\mathbf{K}d^{3}x\Big{)}$
@@
where $\mathbf{h}_{ab}=g_{ab}-\epsilon\mathbf{u}_{a}\mathbf{u}_{b}$ is the induced metric of an arbitrary timelike/spacelike surface (with normal being $u_{a}$) which encloses the boundary, $\mathbf{h}$ is the determinant of the induced metric, $\epsilon=\mathbf{u}^{i}\mathbf{u}_{i}=\pm 1$ depending upon the surface (spacelike/timelike). Also, in the above 30, the well-posed action ($\mathcal{A}_{WP}$) is shown to have consisted of the Einstein-Hilbert action ($\mathcal{A}_{EH}$) along with the GHY boundary term. Furthermore, we use bold text for the quantities of the general surface (timelike or spacelike) to distinguish them from the quantities of the timelike surface which is discussed below. In the above 30, $\boldsymbol{\pi}_{ab}$ can be identified as the gravitational momentum, which is given as
@@
where $\mathbf{K}_{ab}=-\mathbf{h}^{i}_{a}\mathbf{h}^{j}_{b}\nabla_{i}\mathbf{u}_{j}=-\mathbf{h}^{i}_{a}\nabla_{i}\mathbf{u}_{b}$ is the extrinsic curvature of the surface and $\mathbf{K}$ is its trace. The total three-derivative term of 30, $\mathbf{D}_{a}\mathbf{T}^{a}$ (where $\mathbf{T}^{a}=\mathbf{h}^{a}_{i}\mathbf{u}_{j}\delta g^{ij}/16\pi$, and the definition of the three-derivative on the surface is given as $\mathbf{D}_{a}A^{b}=\mathbf{h}_{a}^{i}\mathbf{h}^{b}_{j}\nabla_{i}(\mathbf{h}^{j}_{k}A^{k})$, where $A^{b}$ is an arbitrary vector) can be neglected on the surface and when one extremizes the action.
On a timelike boundary $\mathcal{T}$ (characterised by the spacelike normal $s^{a}$, thereby $\epsilon=+1$, and the induced metric $h_{ab}=g_{ab}-s_{a}s_{b}$), the onshell variation of the well-posed Lagrangian ($\mathcal{L}_{WP}$, where $\mathcal{A}_{WP}=\int_{\mathcal{V}}\sqrt{-g}L_{WP}~{}d^{4}x$) with respect to the induced metric is given as
$\frac{\delta(\sqrt{-g}L_{WP})}{\delta h^{ab}}\Big{|}_{\mathcal{T}}=\pi_{ab}\Big{|}_{\mathcal{T}}=-\frac{T_{ab}^{(BY)}}{2}~{}.$ (32)
Thus, $T_{ab}^{(BY)}$ resembles to the expression of the energy-momentum tensor of the external matter field i.e.
$T_{ab}^{(BY)}=-2\frac{\delta(\sqrt{-g}L_{WP})}{\delta h^{ab}}\Big{|}_{\mathcal{T}}=\frac{1}{8\pi}\Big{(}K_{ab}-Kh_{ab}\Big{)}~{}.$ (33)
Hence, $T_{ab}^{(BY)}$ is known as the surface energy-momentum tensor A.K.A. Brown-York (energy-momentum) tensor. It can be shown that the three-derivative of the BY tensor is given as
$D_{a}T^{ab}_{(BY)}=-T^{ac}_{(ext)}s_{a}h^{b}_{c}~{},$ (34)
where $T^{ab}_{(ext)}$ corresponds to the energy-momentum tensor of the external matter source. In the absence of the external matter, one can obtain $D_{a}T^{ab}_{(BY)}=0$, which resembles to the conservation of the external energy-momentum tensor $\nabla_{a}T^{ab}_{(ext)}=0$ It can be shown that the Brown-York tensor, as defined in 33, is related to several quasi-local parameters of the black holes such as quasi-local energy (popularly known as the Brown-York energy), quasi-local mass (Brown-York mass), angular momentum density etc. (for details see [55, 56]).

### v2

@@ 2.1 Fluid equations
In the discussion of (relativistic) fluid mechanics, the constitutive relations are obtained from an energy-momentum tensor (EMT) and baryon current. The EMT $T^{ab}$ encapsulates the density and flux of energy and momentum in spacetime, providing a comprehensive description of the distribution and flow of energy and momentum within a fluid. The components of the energy-momentum tensor $T^{00}_{\rm(fluid)}$, $T^{0\alpha}_{\rm(fluid)}$ and $T^{\mu\nu}_{\rm(fluid)}$ represent energy density, momentum density and stress respectively. Similarly, the baryon current is also conserved which is expressed via the continuity equation. In relativistic hydrodynamics, there are five independent parameters, three components of fluid velocity (relativistic four components of velocity are constraint by $v_{a}v^{a}=-1$), pressure, and (rest) mass density, that are required to describe fluid completely. In order to uniquely determine these dynamical variables, we require the relativistic Euler equations, derived from the conservation of the energy-momentum tensor and the continuity equation for the conserved rest-mass density. The conservation of the EMT is given by,
$\nabla_{a}T^{ab}_{\rm(fluid)}=0~.$ (1)
The conservation of the EMT (Eq. () yields four equations that describe the evolution of energy and momentum, which are coupled to the pressure, velocity, and density. This can be shown as follows. It is known that if one takes a timelike projection of the conservation law Eq. (, i.e.
$v_{b}\nabla_{a}T^{ab}_{\rm(fluid)}=0~,$ (2)
@@
it yields the energy conservation relation of the fluid, where $v_{b}$ is the four-velocity of the fluid. On the other hand, if one considers the spacelike projection of Eq. (, i.e.
$\Delta^{c}_{b}~\nabla_{a}T^{ab}_{\rm(fluid)}=0~,$ (3)
@@
provides the momentum conservation of the fluid, where $\Delta^{a}_{b}=\delta^{a}_{b}+v^{a}v_{b}$ is the projection tensor, which is orthogonal to the four-velocity $v_{a}$. Note that the above Eq. ( is a set of three dynamical equations. In addition to that, we have the baryon conservation equation for the baryon current $J^{a}$ given as,
$\nabla_{a}J^{a}=0~.$ (4)
To close the system, we require a thermodynamic equation of state. Thus, Eq. (, Eq. ( and Eq. ( together with the equation of state make a set of five closed equations that are required to determine the five unknown quantities stated earlier. Translating to the non-relativistic case, the solution of the five dynamical variables are enabled via the three components of the Navier-Stokes equation, the mass continuity equation for the fluid density and an equation of state (to close the system) that typically relates pressure to fluid density and internal energy (or temperature). When the equation of state depends on temperature, then an additional equation for heat flow has to be supplemented.
@@ 2.2 A generic null hypersurface
[…]
@@
with $\hat{\rho}$ being a smooth scalar field defined on the null surface $\mathcal{H}$. The negative sign in (Eq. () accounts for the fact the null generators $l^{a}$ to $\mathcal{H}$ are future-directed by an appropriate choice of the scalar function $u(x^{a})$. On account of the null nature of the normal to $\mathcal{H}$, $\bm{{l}}$ cannot be provided a unique normalization since $\bm{{l}}\cdot\bm{{l}}=0$. The fact that we have considered only a single null surface, which is the support of the null generator $\bm{{l}}$, we cannot take notions of derivatives of quantities away from the null surface $\mathcal{H}$. To circumvent this problem, following Carter [52], we consider not just a single null surface but rather a family of them. That is, we foliate the spacetime $(\mathcal{M},\textbf{g},\bm{\nabla})$ in the neighborhood of $\mathcal{H}$ by a stack of integrable null surfaces defined by $u(x^{a})=c$, where $c$ is a constant. So we have a family of integrable null surfaces $\mathcal{H}_{u}$ out of which $\mathcal{H}_{u=0}=\mathcal{H}$ is our chosen integrable null hypersurface. This construction then extends the support of the null generator $\bm{{l}}$ to not just over $\mathcal{H}$ but rather over a finite region in the spacetime. This then facilitates the notion of taking derivatives of quantities away from the null surface. It is worth mentioning that such a foliation of the ambient spacetime in the neighborhood of $\mathcal{H}$ is non-unique. However, all geometric quantities that will be described for the null surface will be independent of any such foliation that has been introduced. The integrability of the null surfaces implies that it satisfies the Frobenius identity,
@@
The fact, that the null normal lies on $\mathcal{H}$ itself, does not allow the construction of a projection tensor (that can be created out from the ambient spacetime metric and the null normal) onto the null surface. For that, we need to have the notion of a vector field that is transverse to the null hypersurface. But before introducing such a transverse vector field, let us consider a $3+1$ foliation of the null family $u(x^{a})=c$ by a stack of spacelike surfaces $\Sigma_{t}$ defined by the scalar function $t(x^{a})=\text{constant}$. The spacelike surfaces are assumed not to intersect with each other. Hence, a local neighborhood of the spacetime manifold in the vicinity of $\mathcal{H}$ can be coordinatized by $(t,x^{1},x^{2},x^{3})$, with $x^{\mu}=(x^{1},x^{2},x^{3})$ being the spatial coordinates on a given $\Sigma_{t}$. The time evolution vector field $\bm{t}=\bm{\partial_{t}}$ that connects the same spatial points $x^{\mu}$ of the neighboring slices $\Sigma_{t}$ satisfies the condition $t^{a}\partial_{a}t=1$. The cuts/intersections of the null surface $\mathcal{H}$ with the spacelike familty $\Sigma_{t}$ is defined as $\mathcal{B}^{\rm(null)}\equiv\mathcal{H}\cap\Sigma_{t}$ which form a $2$-dimensional submanifold of $\mathcal{H}$. With this, we introduce a unique notion of a transverse auxiliary null vector field $\bm{{k}}$, defined as,
@@
where $\{\bm{e_{{A}}}\}$ are the set of two spacelike basis vectors on $\mathcal{B}^{\rm(null)}$. The null generators $\bm{{l}}$ provide a notion of outgoing null vector field whereas the auxiliary null vector field $\bm{{k}}$ is ingoing w.r.t the codimension two submanifolds $\mathcal{B}^{\rm(null)}$. With the null generators and the auxiliary null vector field, one can define two projection tensors
@@
On the contrary, $q^{a}_{b}$ is the projection tensor onto the two-surface $\mathcal{B}^{\rm(null)}$ upon which both $l^{a}$ and $k^{a}$ are the normals (i.e. cross-section of $\mathcal{H}$). The properties of $q^{a}_{b}$ are as follows:
@@
Unlike $\Pi^{a}_{~b}$, $q^{a}_{b}$ is the induced metric on $\mathcal{B}^{\rm(null)}$. The intrinsic geometry of the null surface is provided by the first fundamental form, which is the metric $\bm{q}$ induced from the ambient metric $\bm{g}$ from the spacetime manifold $(\mathcal{M},\textbf{g},\bm{\nabla})$. The codimension of two surfaces $\mathcal{B}^{\rm(null)}$ forms a submanifold $(\mathcal{B}^{\rm(null)},\bm{q},\bm{~{}^{(2)}D})$, equipped with the unique metric compatible torsion-free connection $\bm{~{}^{(2)}D}$ satisfying $^{(2)}D_{A}q_{BC}=0$. To completely characterize the extrinsic geometry of the null surface, we need the triplet $(\theta_{ab},\Omega_{a},\kappa)$, where $\theta_{ab}$ is the second fundamental form of $\mathcal{H}$ defined as [25],
@@
In principle, the Hajicek one form is the projection of something called the rotation one-form $\omega_{a}$ onto the submanifold $(\mathcal{B}^{\rm(null)},\bm{q},\bm{~{}^{(2)}D})$, defined as [25],
@@
$q_{a}^{~i}q_{b}^{~j}\Big(\pounds_{\bm{{l}}}\theta_{ab}\Big)=\kappa\theta_{ab}+\theta_{ai}\theta^{i}_{~b}-q_{a}^{~i}q_{b}^{~j}C_{minj}l^{m}l^{\rm N}-\frac{1}{2}q_{ab}R_{ij}l^{i}l^{j}~,$ (18)
where $C_{abcd}$ is the Weyl tensor for $(\mathcal{M},\textbf{g},\bm{\nabla})$. Taking the trace of (Eq. () gives us the celebrated null Raychaudhuri equation (NRE),
@@
and restricting to the spatial coordinates of $(\mathcal{B}^{\rm(null)},\bm{q},\bm{~{}^{(2)}D})$, the above (Eq. () gives us the Hajicek equation [55].
Before ending this geometrical description of the null surface, it is important to mention that we are obviously considering a null surface at finite distances of the spacetime and not null infinity at the conformal boundary of the spacetime conformal completion.
@@ 3.1 BY tensor: laying the path for null counterpart
$\delta\mathcal{A}_{\rm WP}=\delta\mathcal{A}_{\rm EH}+\delta\mathcal{A}_{\rm GHY}=\delta\Big(\frac{1}{16\pi}\int_{\mathcal{V}}\sqrt{-g}R\ d^{4}x-\frac{\epsilon}{8\pi}\int_{\partial\mathcal{V}}\sqrt{\mathbf{h}}\mathbf{K}d^{3}x\Big)$
@@
where $\mathbf{h}_{ab}=g_{ab}-\epsilon\mathbf{u}_{a}\mathbf{u}_{b}$ is the induced metric of an arbitrary timelike/spacelike surface (with normal being $u_{a}$) which encloses the boundary, $\mathbf{h}$ is the determinant of the induced metric, $\epsilon=\mathbf{u}^{i}\mathbf{u}_{i}=\pm 1$ depending upon the surface (spacelike/timelike). Also, in the above Eq. (, the well-posed action ($\mathcal{A}_{\rm WP}$) is shown to have consisted of the Einstein-Hilbert action ($\mathcal{A}_{\rm EH}$) along with the GHY boundary term. Furthermore, we use bold text for the quantities of the general surface (timelike or spacelike) to distinguish them from the quantities of the timelike surface which is discussed below. In the above Eq. (, $\boldsymbol{\pi}_{ab}$ can be identified as the gravitational momentum, which is given as
@@
where $\mathbf{K}_{ab}=-\mathbf{h}^{i}_{a}\mathbf{h}^{j}_{b}\nabla_{i}\mathbf{u}_{j}=-\mathbf{h}^{i}_{a}\nabla_{i}\mathbf{u}_{b}$ is the extrinsic curvature of the surface and $\mathbf{K}$ is its trace. The total three-derivative term of Eq. (, $\mathbf{D}_{a}\mathbf{T}^{a}$ (where $\mathbf{T}^{a}=\mathbf{h}^{a}_{i}\mathbf{u}_{j}\delta g^{ij}/16\pi$, and the definition of the three-derivative on the surface is given as $\mathbf{D}_{a}A^{b}=\mathbf{h}_{a}^{i}\mathbf{h}^{b}_{j}\nabla_{i}(\mathbf{h}^{j}_{k}A^{k})$, where $A^{b}$ is an arbitrary vector) can be neglected on the surface and when one extremizes the action.
On a timelike boundary $\mathcal{T}$ (characterised by the spacelike normal $s^{a}$, thereby $\epsilon=+1$, and the induced metric $h_{ab}=g_{ab}-s_{a}s_{b}$), the onshell variation of the well-posed Lagrangian ($\mathcal{L}_{\rm WP}$, where $\mathcal{A}_{\rm WP}=\int_{\mathcal{V}}\sqrt{-g}L_{\rm WP}~d^{4}x$) with respect to the induced metric is given as
$\frac{\delta(\sqrt{-g}L_{\rm WP})}{\delta h^{ab}}\Big|_{\mathcal{T}}=\pi_{ab}\Big|_{\mathcal{T}}=-\frac{T_{ab}^{\rm(BY)}}{2}~.$ (24)
Thus, $T_{ab}^{\rm(BY)}$ resembles to the expression of the energy-momentum tensor of the external matter field i.e.
$T_{ab}^{\rm(BY)}=-2\frac{\delta(\sqrt{-g}L_{\rm WP})}{\delta h^{ab}}\Big|_{\mathcal{T}}=\frac{1}{8\pi}\Big(K_{ab}-Kh_{ab}\Big)~.$ (25)
Hence, $T_{ab}^{\rm(BY)}$ is known as the surface energy-momentum tensor A.K.A. Brown-York (energy-momentum) tensor. It can be shown that the three-derivative of the BY tensor is given as
$D_{a}T^{ab}_{\rm(BY)}=-T^{ac}_{\rm(ext)}s_{a}h^{b}_{c}~,$ (26)
where $T^{ab}_{\rm(ext)}$ corresponds to the energy-momentum tensor of the external matter source. In the absence of the external matter, one can obtain $D_{a}T^{ab}_{\rm(BY)}=0$, which resembles to the conservation of the external energy-momentum tensor $\nabla_{a}T^{ab}_{\rm(ext)}=0$ . It can be shown that the Brown-York tensor, as defined in Eq. (, is related to several quasi-local parameters of the black holes such as quasi-local energy (popularly known as the Brown-York energy), quasi-local mass (Brown-York mass), angular momentum density etc. (for details see [60, 61]).